# Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding

### v2

@@ Abstract
The primary aim of Knowledge Graph Embeddings (KGE) is to learn low-dimensional representations of entities and relations for predicting missing facts. Although rotation-based methods like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) perform well in KGE, they face two challenges: limited model flexibility requiring proportional increases in relation size with entity dimension, and difficulties in generalizing the model for higher-dimensional rotations. To address these issues, we introduce OrthogonalE, a novel KGE model employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. This approach not only enhances the generality and flexibility of KGE models but also captures several relation patterns that rotation-based methods can identify. Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters. Our code is available at OrthogonalE
@@ 1 Introduction
Recently, rotation-based KGE methods have achieved notable success in the field. For instance, RotatE Sun et al. (2019) conceptualizes relations as 2D rotations while QuatE Zhang et al. (2019) employs 3D rotations to obtain a more expressive model than RotatE. Essentially, as illustrated in Fig. 1 , both operate by multiplying the relation matrix $\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}$ composed of Block-Diagonal Rotation matrix $\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ (RotatE: $\mathbb{R}^{2\times 2}$, QuatE: $\mathbb{R}^{3\times 3}$) with the head entity vector $\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}$.
Figure 1: Fundamental operations ($\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{h}^{1}\approx\mathbf{e}_{t}^{1}$) and inherent challenges of rotation-based KGE models. Rotation-based methods require increasing relation parameters for adequate entity representation and struggle with researching higher-dimensional rotation embeddings ($d>3$) due to their complexity. OrthogonalE, depicted in Fig. 2, efficiently resolves these challenges.
Nevertheless, these approaches encounter two primary issues, as illustrated in Fig. 1. First, the model’s lack of flexibility necessitates increasing the overall relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{(n+l)\times(n+l)}$) to meet entity dimension requirements ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{n+l}$) for better represent entities. For example, when the entity vector changes ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{100}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{1000}$), the parameter increase is 900, but the corresponding change in the relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{100\times 100}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{1000\times 1000}$) results in a parameter increase of 990,000. This substantial increase leads to redundancy and inefficiency in representing relations.
Second, exploring high-dimensional rotational KGE models is challenging due to the significant computational demands and complexity of rotations in higher dimensions ($\mathbf{B}_{i}:\mathbb{R}^{2\times 2},\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{d\times d},d>3$), such as SO(4), SO(5), and SO(10). This restricts the development of more generalized and higher-dimensional rotation KGE approaches.
@@
Table 1: Explanation of notations. Within the table, $\mathbf{e}_{v}$ includes the head $\mathbf{e}_{h}$ and tail $\mathbf{e}_{t}$ entity vectors as used in traditional KGE methods, whereas $\mathbf{e}_{V}$ consists of the head $\mathbf{e}_{H}$ and tail $\mathbf{e}_{T}$ entity matrix representations in our OrthogonalE approach. Furthermore, ’rep’ in the table denotes representation.
To overcome these two issues, we propose a highly general and flexible KGE model named OrthogonalE as shown in Fig. 2, and detail of the notations has been shown in Table. 1. Firstly, by transforming entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ for better represent entities, we control the entity dimension through variable $m$, avoiding unnecessary expansion of the relation size. Corresponding to the above example, we can maintain relation size ($\mathbf{e}_{R}\in\mathbb{R}^{100\times 100}$) and only modify entity matrix size ($\mathbf{e}_{H}^{1}\in\mathbb{R}^{100\times 1}\rightarrow\mathbf{e}_{H}^{2}\in\mathbb{R}^{100\times 10},m:1\rightarrow 10$) to meet the requirements of entity representation. Secondly, leveraging the concept that rotation matrices are orthogonal, we replace rotation matrices $\mathbf{B}_{i}$ with orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ of adaptable dimensions $d$, facilitating the exploration of higher-dimensional block-diagonal orthogonal matrix models. Lastly, for effective optimization, we employ Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and Stochastic Gradient Descent (SGD) for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
@@ 2 Related Work
In conclusion, considering the two major disadvantages of rotation-based methods mentioned in the Introduction 1, there is a need to refine our model to make it more general and flexible.
@@
In optimization on the orthogonal manifold, transitioning from $X^{t}$ to $X^{t+1}$ while remaining on the manifold necessitates a method known as retraction Absil and Malick (2012). Prior research has effectively adapted several standard Euclidean function minimization algorithms to Riemannian manifolds. Notable examples include gradient descent (Absil et al. (2008); Zhang and Sra (2016)), second-order quasi-Newton methods (Absil et al. (2007); Qi et al. (2010)), and stochastic approaches Bonnabel (2013), crucial in deep neural network training.
Meanwhile, we often use Riemannian Optimization for the orthogonal manifold, which has also progressed in deep learning, especially in CNNs and RNNs. Cho and Lee (2017) innovatively substituted CNN’s Batch Normalization layers with Riemannian optimization on the Grassmann manifold for parameter normalization. Additionally, significant strides in stabilizing RNN training have been made by Vorontsov et al. (2017), Wisdom et al. (2016), Lezcano-Casado and Martınez-Rubio (2019), and Helfrich et al. (2018), through the application of Riemannian optimization to unitary matrices.
As this paper primarily focuses on KGE, we do not delve deeply into Riemannian optimization. Instead, we utilize the retraction with exponential map method for iterative optimization, sourced from the Geoopt Kochurov et al. (2020).
@@ 3 Problem Formulation and Background
We present the KGE problem and describe Optimization on the orthogonal manifold before our approach part.
@@ 3.1 Knowledge Graph Embedding
$d^{E}\left(\mathbf{x},\mathbf{y}\right)=\left\|\mathbf{x}-\mathbf{y}\right\|$
@@ 3.2 Optimization on the orthogonal manifold
Figure 3: Abstract representation of Riemannian gradient descent iteration on orthogonal manifold
@@
$\min_{X\in\mathcal{O}_{d}}f(X),$ (1)
Here, $f$ is a differentiable function mapping elements of $\mathbb{R}^{d\times d}$ to $\mathbb{R}$, and the orthogonal manifold $\mathcal{O}_{d}$ is defined as $\mathcal{O}_{d}\triangleq\left\{X\in\mathbb{R}^{d\times d}\mid XX^{\top}=I_{d}\right\}$. Moreover, the tangent space at $X$, denoted by $\mathcal{T}_{X}$, is the set $\mathcal{T}_{X}=\left\{\xi\in\mathbb{R}^{d\times d}\mid\xi X^{\top}+X\xi^{\top}=0\right\}$.
To address problem 1 more efficiently, recent studies suggest optimization on the orthogonal manifold with retractions as an effective approach Ablin and Peyré (2022). In this work, we primarily employ the retraction with exponential map method for iterative optimization, as illustrated in Fig. 3. The key iteration formula for this method is:
@@
The exponential map allows movement in a specified direction on the manifold. Starting from $X$ with initial velocity $\xi$, the exponential map for the orthogonal matrices manifold is represented by Massart and Abrol (2022):
@@
$\text{Grad}f(X)=P_{\mathcal{T}_{X}}(\bigtriangledown f(X)),$
Where $\bigtriangledown f(X)$ is Euclidean gradient of $f(X)$, and the calculation formula for $P_{\mathcal{T}_{X}}(\cdot)$ is:
$P_{\mathcal{T}_{X}}(Y)=X(\frac{X^{\top}Y-Y^{\top}X}{2}),Y\in\mathbb{R}^{d\times d}$
@@ 4.1 Orthogonal Matrices for Relations
QuatE $n$ $n$ $n*m$
OrthogonalE($d$) $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$
@@
QuatE $n$ $n$ $n*m$
OrthogonalE($d$) $n*m$ $n$ $n*m$
@@
Table 2: The parameter calculations for the KGE models. For all models, the relation matrix size is $n$. The block-diagonal matrix size is 2 for RotatE, 3 for QuatE, and $d$ for OrthogonalE($d\times d$), with an entity matrix column size of $m$ for OrthogonalE. In the table, "Normal" represents the standard parameter calculation, "$m=1$" constrains the column size of the entity matrix to 1 to explore the impact of block-diagonal orthogonal matrices on the model, as analyzed in section 5.2.2. "Uniform Entity" ensures that the entity dimensions are consistent across all models to demonstrate the parameter savings in the relation matrix when using the entity matrix in OrthogonalE, as discussed in section 5.2.4.
To address the challenge of exploring high-dimensional rotational KGE models mentioned in the introduction. We exploit the orthogonality of rotation matrices, substituting rotation matrices ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$) with orthogonal matrices ($\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$) of corresponding dimensions $d$. Consequently, our relation embedding ($\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$) are composed of $n/d$ block-diagonal orthogonal matrices $\mathbf{X}_{i}$ as illustrated in Fig. 2:
$\mathbf{e}_{R}=diag(\mathbf{X}_{1},\mathbf{X}_{2},...,\mathbf{X}_{n/d})$ (3)
Where the number of relation parameters is $\frac{d(d-1)}{2}*\frac{n}{d}=\frac{(d-1)n}{2}$, which shown in Table. 2. And this aspect allows OrthogonalE to gain generality, adapting to datasets with diverse complexities by modifying the block-diagonal matrices’ dimension $d$. Additionally, the employed relation structure facilitates the model’s capability to concurrently capture Symmetry, Antisymmetry, Inversion, and Non-commutative Composition relation patterns, as substantiated in Appendix A.2, and detailed introduction of relation patterns refer to Appendix A.4.
4.2 Matrices Representation for Entities
Inspired by Miyato et al. (2022), transforms vector embeddings into matrix embeddings to enhance embedding effectiveness. In our work, to enhance OrthogonalE’s flexibility, we aim to regulate entity dimension using variable $m$ and transform entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ as shown in Fig. 2, thus preventing unnecessary expansion of the relation size. This part allows OrthogonalE to acquire flexibility, adapting to diverse datasets with varying relation and entity parameters, rather than indiscriminately increase both. And the number of entity parameters is $n*m$.
@@ 4.3 Scoring function and Loss
$s(h,r,t)=-d^{E}\left(\mathbf{e}_{R}\cdot\mathbf{e}_{H},\mathbf{e}_{T}\right)+\!b_{h}\!+\!b_{t}$ (4)
Here, $b_{v}(v\in\mathcal{V})$ denotes the entity bias, incorporated as a margin in the scoring function, following methodologies from Tifrea et al. (2018); Balazevic et al. (2019). Furthermore, we opt for uniform selection of negative samples for a given triple $(h,r,t)$ by altering the tail entity, rather than employing alternative negative sampling techniques. The loss function defined as follows:
@@
$y_{t^{\prime}}=\left\{\begin{array}[]{l}-1,\text{ if }t^{\prime}=t\\ 1,\text{ otherwise }\end{array}\right.$ (7)
@@
Table 3: Link prediction accuracy results of two datasets, Bold indicates the best score, and underline represents the second-best score. For a fair comparison, we standardized $m$ at 1 for Gram-Schmidt and all OrthogonalE sizes. The entity dimension for WN18RR was set to approximately 500 (e.g. 501 for 3$\times$3 blocks to ensure experimental feasibility), and around 1000 for FB15K-237. [$\diamondsuit$]: Results are sourced from Zhang et al. (2019). For a fair comparison, the results of RotatE, QuatE, HopfE, and DensE are reported without self-adversarial negative sampling, type constraints, semantics, or reciprocals. More baseline results have been shown in Appendix A.5.

### v3

@@ Abstract
The primary aim of Knowledge Graph Embeddings (KGE) is to learn low-dimensional representations of entities and relations for predicting missing facts. Although rotation-based methods like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) perform well in KGE, they face two challenges: limited model flexibility requiring proportional increases in relation size with entity dimension, and difficulties in generalizing the model for higher-dimensional rotations. To address these issues, we introduce OrthogonalE, a novel KGE model employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. This approach not only enhances the generality and flexibility of KGE models but also captures several relation patterns that rotation-based methods can identify. Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters.
@@
Our code is available at https://github.com/YihuaZhu111/OrthogonalE.
@@ 1 Introduction
Figure 1: Fundamental operations ($\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{h}^{1}\approx\mathbf{e}_{t}^{1}$) and inherent challenges of rotation-based KGE models. Rotation-based methods require increasing relation parameters for adequate entity representation (lack of flexibility) and struggle with researching higher-dimensional rotation embeddings ($d>3$) due to their complexity (lack of generality). OrthogonalE, depicted in Fig. 2, efficiently resolves these challenges.
@@
Recently, rotation-based KGE like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) methods have achieved notable success in the field. RotatE uses the Hadamard product to multiply the real and imaginary components of the head entity embedding by the angle-based relation embedding, resulting in a 2D rotation effect within each unit. Each unit consists of two elements representing the real and imaginary components. For example, an entity embedding with 500 dimensions consists of 250 units. The resulting vectors in each unit are then concatenated to form the tail entity embedding. The essence of this operation is the matrix multiplication between a 2D rotation matrix and the entity vector, followed by concatenation as well, as illustrated in Fig. 1, multiplying the relation matrix $\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}$ composed of the block-diagonal Rotation matrix $\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ (RotatE: $\mathbb{R}^{2\times 2}$) with the head entity vector $\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}$. Here, we treat the concatenation operation as a block-diagonal arrangement of each 2D rotation matrix. Similarly, QuatE extends RotatE by using quaternions, which consist of one real and three imaginary components, achieving a 3D rotation effect through the multiplication of a 3D rotation matrix with the entity vector. For QuatE, we replace the 2D rotation matrix $\mathbb{R}^{2\times 2}$ with a 3D rotation matrix $\mathbb{R}^{3\times 3}$ as shown in Fig. 1.
However, these approaches face two primary issues: lack of flexibility and generality, as illustrated in Fig. 1. First, the model’s lack of flexibility necessitates increasing the size of the overall relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{(n+l)\times(n+l)}$) to meet entity dimension requirements ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{n+l}$) for better entity representation. For example, when the entity vector changes ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{100}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{1000}$) to improve representation, the parameter increase is 900. However, the corresponding change in the relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{100\times 100}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{1000\times 1000}$) results in a parameter increase of 990,000. This substantial increase leads to redundancy and inefficiency in representing relations.
[…]
@@
Table 1: Notation summary. Within the table, $\mathbf{e}_{v}$ includes the head $\mathbf{e}_{h}$ and tail $\mathbf{e}_{t}$ entity vectors as used in traditional KGE methods, whereas $\mathbf{e}_{V}$ consists of the head $\mathbf{e}_{H}$ and tail $\mathbf{e}_{T}$ entity matrix representations in our OrthogonalE approach. Furthermore, ‘rep’ in the table denotes representation.
[…]
@@ 2 Related Work
In addition, our research builds upon existing work in the field. For instance, OTE Tang et al. (2019) uses a compact block-diagonal orthogonal matrix, similar to our approach, to maintain RotatE’s relation patterns and complex relations. Yet, our OrthogonalE model improves stability and performance with Riemannian Optimization, outperforming OTE’s Gram-Schmidt process. GoldE’s Li et al. (2024) universal orthogonal parameterization, derived from Householder reflections, offers theoretical guarantees for dimensionality and geometry. However, our focus is on enhancing the generality and flexibility of the KGE models, not just on capturing knowledge graph patterns and heterogeneity.
In conclusion, considering the two major disadvantages of rotation-based methods mentioned in Section 1, we need to refine our model to make it more general and flexible.
@@
In optimization on the orthogonal manifold, transitioning from $X^{t}$ to $X^{t+1}$ while remaining on the manifold necessitates a method known as retraction Absil and Malick (2012). Prior research has effectively adapted several standard Euclidean function minimization algorithms to Riemannian manifolds. Notable examples include gradient descent Absil et al. (2008); Zhang and Sra (2016), second-order quasi-Newton methods Absil et al. (2007); Qi et al. (2010), and stochastic approaches Bonnabel (2013), crucial in deep neural network training.
Meanwhile, Riemannian optimization is often used for the orthogonal manifolds, and has made significant progress in deep learning, especially in CNNs and RNNs. Cho and Lee (2017) innovatively substituted CNN’s Batch Normalization layers with Riemannian optimization on the Grassmann manifold for parameter normalization. Additionally, significant strides in stabilizing RNN training have been made by Vorontsov et al. (2017); Wisdom et al. (2016); Lezcano-Casado and Martınez-Rubio (2019); Helfrich et al. (2018), through the application of Riemannian optimization to unitary matrices.
As this paper primarily focuses on KGE, we do not delve deeply into Riemannian optimization. Instead, we utilize retraction with the exponential map for iterative optimization, as implemented in the Geoopt library Kochurov et al. (2020).
@@ 3 Problem Formulation and Background
Before presenting our approach, we introduce the KGE problem and provide an overview of optimization on orthogonal manifolds.
@@ 3.1 Knowledge Graph Embedding
$d^{E}\left(\mathbf{x},\mathbf{y}\right)=\left\|\mathbf{x}-\mathbf{y}\right\|.$
@@ 3.2 Optimization on the orthogonal manifold
Figure 3: Illustration of Riemannian gradient descent iteration on an orthogonal manifold.
@@
$\min_{X\in\mathcal{O}_{d}}f(X).$ (1)
Here, $f$ is a differentiable function mapping elements of $\mathbb{R}^{d\times d}$ to $\mathbb{R}$, and the orthogonal manifold $\mathcal{O}_{d}$, also denoted as $O(d)$, is defined as $\mathcal{O}_{d}\triangleq\left\{X\in\mathbb{R}^{d\times d}\mid XX^{\top}=I_{d}\right\}$. Moreover, the tangent space at $X$, denoted by $\mathcal{T}_{X}$, is the set $\mathcal{T}_{X}=\left\{\xi\in\mathbb{R}^{d\times d}\mid\xi X^{\top}+X\xi^{\top}=0\right\}$.
To address the problem formulated in (1) more efficiently, recent studies suggest optimization of the orthogonal manifold with retractions as an effective approach Ablin and Peyré (2022). In this work, we primarily employ the retraction with exponential map for iterative optimization, as illustrated in Fig. 3. The key iteration formula for this method is:
@@
The exponential map allows movement in a specified direction on the manifold. Starting from $X$ with initial velocity $\xi$, the exponential map on the manifold of orthogonal matrices is given by Massart and Abrol (2022):
@@
$\text{Grad}f(X)=P_{\mathcal{T}_{X}}(\nabla f(X)),$
where $\nabla f(X)$ is Euclidean gradient of $f(X)$, and the calculation formula for $P_{\mathcal{T}_{X}}(\cdot)$ is:
$P_{\mathcal{T}_{X}}(Y)=X(\frac{X^{\top}Y-Y^{\top}X}{2}),Y\in\mathbb{R}^{d\times d}.$
@@ 4.1 Orthogonal Matrices for Relations
QuatE $n$ $n$ $n\times m$
OrthogonalE($d\times d$) $n\times m$ $n$ $n\times m$
@@
QuatE $n$ $n$ $n\times m$
OrthogonalE($d\times d$) $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$
@@
Table 2: The parameter calculations for the KGE models. For all models, the relation matrix size is $n$. The block-diagonal matrix size is 2 for RotatE, 3 for QuatE, and $d$ for OrthogonalE, with an entity matrix column size of $m$ for OrthogonalE. In the table, “Normal” represents the standard parameter calculation, “$m=1$” constrains the column size of the entity matrix to 1 to explore the impact of block-diagonal orthogonal matrices on the model, as analyzed in section 5.2.2. “Fixed $\mathbf{e}_{V}$” ensures that the entity dimensions are consistent across all models to demonstrate the parameter savings in the relation matrix when using the entity matrix in OrthogonalE, as discussed in section 5.2.4.
To address the challenge of generality for exploring high-dimensional rotational KGE models mentioned in Section 1, we exploit the orthogonality of rotation matrices, substituting rotation matrices ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$) with orthogonal matrices ($\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$) of corresponding dimensions $d$. Consequently, our relation embedding ($\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$) is composed of $n/d$ block-diagonal orthogonal matrices $\mathbf{X}_{i}$ as illustrated in Fig. 2:
$\mathbf{e}_{R}=\textrm{diag}(\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{n/d}),$ (3)
where the number of relation parameters is $\tfrac{d(d-1)}{2}\times\tfrac{n}{d}=\tfrac{(d-1)n}{2}$ as shown in Table 2. This structure allows OrthogonalE to achieve generality, adapting to datasets with diverse complexities by adjusting the dimension $d$ of the block-diagonal matrices. Additionally, the employed relation structure facilitates the model’s capability to concurrently capture Symmetry, Antisymmetry, Inversion, and Non-commutative Composition relation patterns, as substantiated in Appendix A.2, and detailed introduction of relation patterns refer to Appendix A.4.
4.2 Matrix Representation for Entities
Inspired by Miyato et al. (2022), transforms vector embeddings into matrix embeddings to improve embedding effectiveness. In our work, to enhance OrthogonalE’s flexibility, we aim to regulate entity dimension using variable $m$ and transform entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ as shown in Fig. 2, thus preventing unnecessary expansion of the relation size as shown in Fig. 1. This part allows OrthogonalE to acquire flexibility, adapting to diverse datasets with varying relation and entity parameters, rather than indiscriminately increase both. And the number of entity parameters is $n\times m$.
@@ 4.3 Scoring function and Loss
$s(h,r,t)=-d^{E}\left(\mathbf{e}_{R}\cdot\mathbf{e}_{H},\mathbf{e}_{T}\right)+\!b_{h}\!+\!b_{t}.$ (4)
Here, $b_{v}\,(v\in\mathcal{V})$ denotes the entity bias, incorporated as a margin in the scoring function, following methodologies from Tifrea et al. (2018); Balazevic et al. (2019). Furthermore, we opt for a uniform selection of negative samples for a given triple $(h,r,t)$ by altering the tail entity, rather than employing alternative negative sampling techniques. The loss function defined as follows:
@@
$y_{t^{\prime}}=\left\{\begin{array}[]{l}-1,\text{ if }t^{\prime}=t\\ 1,\text{ otherwise},\end{array}\right.$ (7)
where $t^{\prime}$ represents sampled tail entities that include both positive and negative samples.
@@
Table 3: Link prediction accuracy results of two datasets, Bold indicates the best score, and underline represents the second-best score. For a fair comparison, we standardized $m$ at 1 for Gram-Schmidt and all OrthogonalE sizes. The entity dimension for WN18RR was set at approximately 500 (for example, 501 for 3$\times$3 blocks to ensure experimental feasibility) and around 1000 for FB15K-237. [$\diamondsuit$]: The results are sourced from Zhang et al. (2019). For a fair comparison, the results of RotatE, QuatE, HopfE, and DensE are reported without self-adversarial negative sampling, type constraints, semantics, or reciprocals. More baseline results are shown in Appendix A.5.